OpenStudy (unavailabilityy):

FAN + MEDAL! Please describe to me what monomials do. What I mean is, what exactly are the steps to solving a monomial equation? Create your own monomial and then show me the steps to solving it. Thank you!

OpenStudy (clinton):

How do you not know something so simple?

OpenStudy (unavailabilityy):

To me monomial expressions are pretty complicated to learn and so I asked this question. The same goes for polynomials and binomials.

OpenStudy (clinton):

You silly goose, this is child's play.

OpenStudy (518nad):

monomial is a mono term expression, so single term expression

OpenStudy (518nad):

x^2 is a monmial

OpenStudy (518nad):

x^2=4 x=+/- 2

OpenStudy (518nad):

or 5x=10 5x/5=10/5 x=2